OpenStudy (anonymous):

Find the difference. (16x3 + 2x2) - 4x3 =

OpenStudy (anonymous):

\[(16x^3 + 2x^2) - 4x^3=(16-4)x^3 + 2x^2=12x^3+2x^2\]

OpenStudy (jamesj):

48 + 4 - 12 = 40

OpenStudy (anonymous):

40

OpenStudy (anonymous):

\[A. 14x ^{3} ..........B.12x ^{3} - 2x ^{2}.....C.12x^{3}+ 2x ^{2}...D.16x ^{3} - 2x ^{2}\]

OpenStudy (anonymous):

my answer matches C